The function described by H (x) = 2.75x+ 71.48 can be used to predict the height, in centimeters, of a woman whose humerous (the bone from the elbow to the shoulder)is x cm long.

Predict the height of a woman whose humerous is 36 cm long.

Found 2 solutions by stanbon, jojo14344:
Answer by stanbon(57974)

H (x) = 2.75x+ 71.48
-----
H(36) = 2.75*36 + 71.48
H(36) = 170.48
----------------
Cheers,
Stan H.

Answer by jojo14344(1512)



Working Eqn,
H(x) = 2.75x + 71.48


Subst. x= 36cm:

H(x)= 2.75(36) + 71.48 = 99 + 71.48
H = 170.48cm


You want to verify, we'll see the graph.



Thank you,
Jojo